sin(x)^2=2cos(x)+2
1
Expert's answer
2012-02-14T10:24:05-0500
Due to the famous identity cos2x+sin2x=1 therefore sin2x=1-cos2x
So sin2x=1-cos2x =2cos x+2
Let's denote cos x=t. We suppose that -1<=t<=1
So we get the equation:
1-t^2=2t+2
t^2+2t+1=0
(t+1)^2=0
t=-1
So cos x=-1
x=pi+2k pi where k - integer

Also we created video where this particular problem is considered. Please take a look!